A hair salon surveyed 221 customers (105 females and 116 males) to see if they are satisfied with the service. The result is summarized in the following table.

Satisfied Not Satisfied Total
Female 100 5 105
Male 114 2 116
Total 214 7 221

1. If a customer is randomly selected from these 221 people, the probability that he/she is satisfied is ??

(Please round your answer to two decimal places.)

2. If we know the selected customer is a female, then what is the probability that she is satisfied?

(Please round your answer to two decimal positions.)

P(satisfied|female) =  ??

3. How about the probability that a randomly selected customer is a female if we know that the person is satisfied?

(Please round your answer to two decimal positions.)

P(female|satisfied) = ??
Answer #1

Using given summarized table we want to find, the following probabilities,

1) P( Satisfied) = 214 / 221 = 0.9683

=> P(Satisfied) = 0.97

2) P(satisfied | female) = 100 / 105 = 0.9524

=> P(satisfied | female ) = 0.95

3) P(female | satisfied) = 100 / 214 = 0.4673

=> P(female | satisfied) = 0.47